# SONIC: Supersizing Motion Tracking for Natural Humanoid Whole-Body Control

### v1

@@ 3.1 Humanoid Motion Dataset
Our motion dataset is built from a large-scale motion-capture corpus comprising recordings from 170 human subjects. The participants include a balanced mix of male and female actors, with body heights ranging from 145 cm to 199 cm (mean = 174.3 cm, std = 10.9 cm), approximately following a normal distribution. The dataset spans a broad spectrum of everyday human behaviors, including locomotion, daily activities, gesturing, and a diverse set of combat motions with varied stylistic expressions. Clip durations range from 1 to 180 seconds. In total, the collection covers thousands of unique motion behaviors, with most actions performed by multiple subjects across multiple takes, providing rich intra- and inter-subject variation, as can be seen in Fig.˜7. Our in-house dataset contains 700 hours of human motion, resulting in 100+ million frames at 50Hz. The human motion is then retargeted to the humanoid using GMR (Araujo et al., 2025).
@@ 3.2 Universal Humanoid Motion Tracking
We formulate humanoid motion tracking as a Markov Decision Process ${\mathcal{M}}=\langle\mathcal{\bm{S}},\mathcal{\bm{A}},\mathcal{\bm{T}},\mathcal{R},\gamma\rangle$, comprising state space, action space, transition function, reward function, and discount factor $\gamma$. We train the policy using proximal policy optimization (PPO) (Schulman et al., 2017) to maximize the expected cumulative discounted return $\mathbb{E}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\right]$.
@@
States. The state representation $\bm{s}_{t}$} comprises two components: proprioceptive sensing $\bm{s}^{\text{p}}_{t}$} and motion command $\bm{s}^{\text{g}}_{t}$}. Proprioceptive information ${\bm{s}^{\text{p}}_{t}}\triangleq({\bm{{q}}_{t}},{\bm{\dot{q}}_{t}},{\bm{{\omega}}_{t}},\bm{\psi}_{t},\bm{a}_{t-1})$ includes joint pose $\bm{{q}}_{t}$}, joint velocity $\bm{\dot{q}}_{t}$}, root angular velocity $\bm{{\omega}}_{t}$}, gravity vector $\bm{g}_{t}$ in the root frame, and previous action $\bm{a}_{t-1}$. The motion command $\bm{s}^{\text{g}}_{t}$} has three types: robot motion $\bm{g}_{r}$, human motion $\bm{g}_{h}$, or hybrid motion $\bm{g}_{m}$ (combining upper-body keypoints with lower-body robot motions), where we drop the subscript $t$ for brevity. All state quantities are expressed in the robot’s local heading frame to ensure rotation invariance.
@@
Actions. The policy $\pi$ outputs target joint positions $\bm{a}_{t}$} as actions, which are tracked by proportional-derivative (PD) controllers at each joint.
@@
Rewards. Following Liao et al. (2025), we define the reward as $r_{t}=\mathcal{R}({\bm{s}^{\text{p}}_{t}},{\bm{s}^{\text{g}}_{t}})+\mathcal{P}({\bm{s}^{\text{p}}_{t}},\bm{a}_{t})$, combining tracking reward and penalty terms. The tracking term $\mathcal{R}$ minimizes errors in root position, root orientation, body link positions (relative to root), body link orientations (relative to root), body link linear velocities, and body link angular velocities between the robot state $\bm{s}^{\text{p}}_{t}$} and target $\bm{s}^{\text{g}}_{t}$}. The penalty term $\mathcal{P}$ discourages abrupt action changes, joint limit violations, and undesired contacts. Detailed reward design is presented in Table 1.
@@
Domain Randomization. To enhance robustness and generalization across diverse scenarios, we apply systematic domain randomization during training. We randomize physical parameters which include friction coefficients ($\mu_{s}$, $\mu_{d}$), restitution coefficient ($e$), default joint positions ($\bm{q}_{0}$), and base center-of-mass position. We also periodically apply random perturbations to the robot’s root linear and angular velocities to simulate external pushes. Additionally, we apply motion perturbation to the target motion commands $\bm{s}^{\text{g}}_{t}$} during training to improve robustness. All domain randomization parameters are detailed in Table 2.
@@
Reward term Equation Weight
Tracking rewards $\mathcal{R}({\bm{s}^{\text{p}}_{t}},{\bm{s}^{\text{g}}_{t}})$
Penalty terms $\mathcal{P}({\bm{s}^{\text{p}}_{t}},\bm{a}_{t})$
Action rate $r_{\text{act}}(t)=\|\bm{a}_{t}-\bm{a}_{t-1}\|_{2}^{2}$ -0.1
Table 1: Reward design. Orientations are represented as 6D rotations (Zhou et al., 2019). Superscript $g$: goal/target state from $\bm{s}^{\text{g}}_{t}$}; superscript $p$: proprioceptive/current state from $\bm{s}^{\text{p}}_{t}$}; $\mathcal{B}$: set of tracked body links; superscript “rel”: quantities expressed in root frame.
Domain Randomization Sampling Distribution
Physical parameters
Static friction coefficients $\mu_{s}\sim\mathcal{U}[0.3,1.6]$
Dynamic friction coefficients $\mu_{d}\sim\mathcal{U}[0.3,1.2]$
Restitution coefficient $e\sim\mathcal{U}[0,0.5]$
Default joint positions $\bm{q}_{0}\sim\bm{q}_{0}+\mathcal{U}[-0.01,0.01]$
Root velocity perturbations (external pushes)
Push duration $\Delta t\sim\mathcal{U}[1,3]\text{s}$
Target motion perturbations ($\bm{s}^{\text{g}}_{t}$})
Target joint jitter $\Delta\bm{q}_{t}^{g}\sim\mathcal{U}[-0.1,0.1]$
Module Architecture Dims
Network configuration
Quantizer FSQ token dimensions $D_{z}$; quantization levels $L_{z}$
Encoder (g1) MLP hidden=[2048,1024,512,512]=[2048,1024,512,512]
Encoder (teleop) MLP hidden =[2048,1024,512,512]=[2048,1024,512,512]
Encoder (smpl) MLP hidden =[2048,1024,512,512]=[2048,1024,512,512]
Decoder (actions) MLP hidden =[2048,2048,1024,1024,512,512]=[2048,2048,1024,1024,512,512]
Decoder (refs) MLP hidden =[2048,1024,512,512]=[2048,1024,512,512]
Action dimension Diagonal Gaussian $29$
Critic MLP hidden =[2048,2048,1024,1024,512,512]=[2048,2048,1024,1024,512,512]
Motion command
Future frames $F_{r}=F_{h}=F_{m}=10$ frames
Frame interval $\Delta t_{r}=t_{m}=0.1s$, $\Delta t_{h}=0.02s$
Table 3: Universal control policy hyperparameters.
Training hyperparameter Value
Num parallel envs per GPU $4096$
Num steps per env 24
Learning epochs 55
Num mini-batches 44
Discount $\gamma$ $0.99$
GAE $\lambda$ $0.95$
Clip parameter $0.2$
Entropy coefficient $0.013$
Value loss coefficient $1.0$
Actor learning rate $2{\times}10^{-5}$
Critic learning rate $1{\times}10^{-3}$
Max gradient norm $0.1$
Desired KL $0.01$
Adaptive LR min/max $1{\times}10^{-5}$ , $2{\times}10^{-4}]$
Init noise std $0.05$
Actor std clamp min/max $[0.001,0.5]$
Adaptive sampling bin size 1s
Adaptive sampling failure rate cap $\beta=200$
Adaptive sampling blending hyperparameter $\alpha=0.1$
Table 4: Training hyperparameters.
@@
We prepare synchronized motion data across all three command types. Each command type $\bm{g}_{r},\bm{g}_{h},\bm{g}_{m}$ is encoded via its respective encoder and quantized to produce universal tokens $\bm{z}_{r},\bm{z}_{h},\bm{z}_{m}$. For each token, the control decoder $\bm{\mathcal{D}}_{c}$ generates motor commands while the motion decoder $\bm{\mathcal{D}}_{r}$ reconstructs the robot motion command. The total loss comprises:
@@
We employ bin-based adaptive motion sampling when selecting the initial frame for each episode. Specifically, the entire motion dataset is partitioned into fixed-duration bins. For each bin, we calculate the corresponding failure rate $f_{i}$ of the policy. To prevent excessive sampling from bins with exceptionally high failure rates, the failure rate in each bin is capped at $\beta\bar{f}$, where $\beta$ is a hyperparameter and $\bar{f}$ is the average failure rate across all bins. The normalized, capped failure rates are then used to derive preliminary sampling weights $\hat{p}_{i}$ for each bin. The final sampling probability is defined as $p_{i}=\alpha\hat{p}_{i}+(1-\alpha)\frac{1}{N}$, where $\alpha$ is a blending hyperparameter and $N$ is the total number of bins. This formulation balances targeted sampling of challenging bins with uniform coverage. A bin is selected according to $p_{i}$, and the initial frame is then sampled uniformly from within the chosen bin. All models are trained using Isaac Lab (NVIDIA et al., 2025) as the simulation platform. Training hyperparameters are provided in Table 4.
@@ 3.3 Generative Kinematic Motion Planner
Our generative kinematic motion planner is a large-scale latent generative model, trained on the same natural whole-body motion data as the motion tracking policy. At a high level, the planning process is formulated as an autoregressive motion in-betweening generation task. The context keyframes capture historical robot states, such as joint positions and root positions, while target keyframes are either navigation guidance keyframes generated from the user commands such as velocity, direction, and style, or skill-specific targets for actions such as squatting, crawling, boxing, etc.
@@
During training, we sample motion segments of length between 0.8s and 2.4s, extracting the keyframes at both endpoints to serve as the context and target keyframes. Our motion representation is mathematically equivalent to the humanoid pose configuration $q_{t}$ as introduced in Sec.˜3.2. Specifically, we represent kinematic motion using the pelvis-relative joint positions and global joint rotations. During training, we randomly rotate the training samples to enable planning in all initial orientations. Incorporating global rotation instead of local, canonicalized rotation is essential for generating motions such as squatting and crawling, where the notion of heading is ill-defined and greatly hurts the motion planning quality. We refer readers to Meng et al. (2025a) and Meng et al. (2025b) for similar insights and additional discussion.
@@ 3.5 Deployment
The policy loop runs at $50$ Hz. At each cycle, we assemble observations from recent robot state and operator intent, evaluate a learned policy, and synthesize joint-space targets. To ensure crisp, compliant actuation, these targets are streamed by a separate high-rate process at $500$ Hz through the Unitree low-level API, continuously publishing the latest user inputs (target motion command) without blocking the policy loop.
@@
Both the interactive kinematic motion planner and the policy inference execute onboard the Jetson Orin GPU using TensorRT with CUDA Graph acceleration. This configuration yields reliable, low-variance execution with $1\text{--}2$ ms per forward pass for policy and $12$ ms per forward pass for motion generation. The resulting architecture emphasizes timing robustness and smoothness. It produces a stable motion, while decoupled loops isolate sensing and planning from control, stabilizing end-to-end latency. Unless otherwise noted, all reported experiments use this onboard configuration (50/10/500/100 Hz) with timing logs enabled for diagnostics and reproducibility.
Appendix A Acknowledgments
We thank Scott Reed, Wei Liu, You Liang Tan, Avnish Narayan, Fengyuan Hu, Qi Wang, Yuqi Xie, Letian (Max) Fu, Mengda Xu, Davis Rempe, Xue Bin (Jason) Peng, Haotian Zhang, Yifeng Jiang, Anna Minx, John Malaska, Chen Tessler for their thoughtful discussions. We thank Leilee Naderi and Amanpreet Singh for supporting with data collection and providing feedback on teleoperation user experience. We thank Jeremy Chimienti and Tri Cao for their work ensuring robot readiness and conducting the necessary repairs. We also thank Tri Cao, Jazmin Sanchez, Demetria Quijada, and Jesse Yang for their help in filming and editing.
Appendix B Contributors

### v2

@@ 3.1 Humanoid Motion Dataset
Our motion dataset is built from a large-scale motion-capture corpus comprising recordings from 170 human subjects. The participants include a balanced mix of male and female actors, with body heights ranging from 145 cm to 199 cm (mean = 174.3 cm, std = 10.9 cm), approximately following a normal distribution. The dataset spans a broad spectrum of everyday human behaviors, including locomotion, daily activities, gesturing, and a diverse set of combat motions with varied stylistic expressions. Clip durations range from 1 to 180 seconds. In total, the collection covers thousands of unique motion behaviors, with most actions performed by multiple subjects across multiple takes, providing rich intra- and inter-subject variation, as can be seen in Fig.˜7. Our in-house dataset contains 700 hours of human motion, resulting in 100+ million frames at 50Hz. The human motion retargeted is adapted from GMR (Araujo et al., 2025) and PyRoki (Kim* et al., 2025).
@@ 3.2 Universal Humanoid Motion Tracking
We formulate humanoid motion tracking as a Markov Decision Process ${\mathcal{M}}=\langle\mathcal{\bm{S}},\mathcal{\bm{A}},\mathcal{\bm{T}},\mathcal{R},\gamma\rangle$, comprising state space, action space, transition function, reward function, and discount factor $\gamma$. We train the policy using proximal policy optimization (PPO) (Schulman et al., 2017) to maximize the expected cumulative discounted return $\mathbb{E}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\right]$. Our environment design follows the general motion tracking formulation (Luo et al., 2023; Chen et al., 2025; Zhang et al., 2025; He et al., 2024; Liao et al., 2025), and we adapt the well-tuned environmental settings from (Liao et al., 2025) as the basis for scaling up humanoid motion tracking.
@@
States. The state representation $\bm{s}_{t}$} comprises two components: proprioceptive sensing $\bm{s}^{\text{p}}_{t}$} and motion command $\bm{s}^{\text{g}}_{t}$}. Proprioceptive information ${\bm{s}^{\text{p}}_{t}}\triangleq({\bm{{q}}_{t}},{\bm{\dot{q}}_{t}},{\bm{{\omega}}_{t}},\bm{\psi}_{t},\bm{a}_{t-1})$ includes joint pose $\bm{{q}}_{t}$}, joint velocity $\bm{\dot{q}}_{t}$}, root angular velocity $\bm{{\omega}}_{t}$}, gravity vector $\bm{g}_{t}$ in the root frame, and previous action $\bm{a}_{t-1}$. The motion command $\bm{s}^{\text{g}}_{t}$} has three types: robot motion $\bm{g}_{r}$, human motion $\bm{g}_{h}$, or hybrid motion $\bm{g}_{m}$ (combining upper-body keypoints with lower-body robot motions), where we drop the subscript $t$ for brevity. All state quantities are expressed in the robot’s local heading frame to ensure rotation invariance. We use the 6D rotations representation (Zhou et al., 2019) throughout.
@@
Actions. The policy $\pi$ outputs target joint positions $\bm{a}_{t}$} as actions, which are tracked by proportional-derivative (PD) controllers at each joint. For PD gains setting, we follow prior art Raibert and Farshidian (2025); Liao et al. (2025) that proves effective in training high-quality tracking policy.
@@
Rewards. We define the reward as $r_{t}=\mathcal{R}({\bm{s}^{\text{p}}_{t}},{\bm{s}^{\text{g}}_{t}})+\mathcal{P}({\bm{s}^{\text{p}}_{t}},\bm{a}_{t})$, combining tracking reward and penalty terms. The tracking term $\mathcal{R}$ minimizes errors in root position, root orientation, body link positions (relative to the root), body link orientations (relative to the root), body link linear velocities, and body link angular velocities between the robot state $\bm{s}^{\text{p}}_{t}$} and the target $\bm{s}^{\text{g}}_{t}$}. The penalty term $\mathcal{P}$ discourages abrupt action changes, joint limit violations, and undesired contacts. Detailed reward design is presented in Table 3.
@@
Domain Randomization. To enhance robustness and generalization across diverse scenarios, we apply systematic domain randomization during training. We randomize physical parameters, which include friction coefficients ($\mu_{s}$, $\mu_{d}$), the restitution coefficient ($e$), default joint positions ($\bm{q}_{0}$), and the base center-of-mass position. We also periodically apply random perturbations to the robot’s root linear and angular velocities to simulate external pushes. Additionally, we apply motion perturbation to the target motion commands $\bm{s}^{\text{g}}_{t}$} during training to improve robustness. All domain randomization parameters are detailed in Table 4.
@@
We prepare synchronized motion data across all three command types. Each command type $\bm{g}_{r},\bm{g}_{h},\bm{g}_{m}$ is encoded via its respective encoder and quantized to produce universal tokens $\bm{z}_{r},\bm{z}_{h},\bm{z}_{m}$. For each token, the control decoder $\bm{\mathcal{D}}_{c}$ generates motor commands, while the motion decoder $\bm{\mathcal{D}}_{r}$ reconstructs the robot motion command. The total loss comprises:
@@
We employ bin-based adaptive motion sampling when selecting the initial frame for each episode. Specifically, the entire motion dataset is partitioned into fixed-duration bins. For each bin, we calculate the corresponding failure rate $f_{i}$ of the policy. To prevent excessive sampling from bins with exceptionally high failure rates, the failure rate in each bin is capped at $\beta\bar{f}$, where $\beta$ is a hyperparameter and $\bar{f}$ is the average failure rate across all bins. The normalized, capped failure rates are then used to derive preliminary sampling weights $\hat{p}_{i}$ for each bin. The final sampling probability is defined as $p_{i}=\alpha\hat{p}_{i}+(1-\alpha)\frac{1}{N}$, where $\alpha$ is a blending hyperparameter and $N$ is the total number of bins. This formulation balances targeted sampling of challenging bins with uniform coverage. A bin is selected according to $p_{i}$, and the initial frame is then sampled uniformly from within the chosen bin. We train our motion tracker using distributed training powered by Gugger et al. (2022) and von Werra et al. (2020) across multiple compute nodes. All models are trained using Isaac Lab (NVIDIA et al., 2025) as the simulation platform. Training hyperparameters are provided in Table 2.
@@ 3.3 Generative Kinematic Motion Planner
Our generative kinematic motion planner is a large-scale latent generative model, trained on the same natural whole-body motion data as the motion tracking policy. At a high level, the planning process is formulated as an autoregressive motion in-betweening generation task. The context keyframes capture historical robot states, such as joint positions and root positions, while target keyframes are either navigation guidance keyframes generated from user commands such as velocity, direction, and style, or skill-specific targets for actions such as squatting, crawling, boxing, etc.
@@
During training, we sample motion segments of length between 0.8s and 2.4s, extracting the keyframes at both endpoints to serve as the context and target keyframes. Our motion representation is mathematically equivalent to the humanoid pose configuration $q_{t}$ as introduced in Sec.˜3.2. Specifically, we represent kinematic motion using the pelvis-relative joint positions and global joint rotations. During training, we randomly rotate the training samples to enable planning in all initial orientations. Incorporating global rotation instead of local, canonicalized rotation is essential for generating motions such as squatting and crawling, where the notion of heading is ill-defined and greatly affects the quality of motion planning. We refer readers to Meng et al. (2024) and Meng et al. (2025) for similar insights and additional discussion.
@@ 3.5 Deployment
The policy loop runs at $50$ Hz. At each cycle, we assemble observations from the recent robot state and operator intent, evaluate a learned policy, and synthesize joint-space targets. To ensure crisp, compliant actuation, these targets are streamed by a separate high-rate process at $500$ Hz through the Unitree low-level API, continuously publishing the latest user inputs (target motion command) without blocking the policy loop.
@@
Both the interactive kinematic motion planner and the policy inference execute onboard the Jetson Orin GPU using TensorRT with CUDA Graph acceleration. This configuration yields reliable, low-variance execution with $1\text{--}2$ ms per forward pass for policy and $12$ ms per forward pass for motion generation. The resulting architecture emphasizes timing robustness and smoothness. It produces a stable motion, while decoupled loops isolate sensing and planning from control, stabilizing end-to-end latency. Unless otherwise noted, all reported experiments use this onboard configuration (50/10/500/100 Hz) with timing logs enabled for diagnostics.
4 Acknowledgments
We thank Scott Reed, Wei Liu, You Liang Tan, Avnish Narayan, Fengyuan Hu, Yuqi Xie, Letian (Max) Fu, Mengda Xu, Davis Rempe, Xue Bin (Jason) Peng, Haotian Zhang, Yifeng Jiang, Anna Minx, John Malaska, Chen Tessler, Soha Pouya, Kaushil Prakashbhai Kundalia, Huihua Zhao, Xiaowei Jiang, Olivier Dionne, Michael De Ruyter, Michael Buttner, Qi Wang, Yeongho Seol, Mathis Petrovich, Sanja Fidler, Kaifeng Zhao, Spencer Huang, Gavriel State, Yurong "Kelly" Guo for their thoughtful discussions and help. We thank Leilee Naderi and Amanpreet Singh for supporting with data collection and providing feedback on teleoperation user experience. We thank Jeremy Chimienti and Tri Cao for their work ensuring robot readiness and conducting the necessary repairs. We also thank Tri Cao, Jazmin Sanchez, Demetria Quijada, and Jesse Yang for their help in filming and editing.
5 Contributors
@@
Appendix A Supplementary Materials
S1. Supplementary Video Descriptions
The project site hosts a collection of high-resolution videos that complement the results presented in the main paper. Each video demonstrates SONIC ’s large-scale motion tracking, interactive control, multi-modal interfaces, and VLA integrations on a Unitree G1. Below we provide detailed descriptions of each video category.
S1.1 Teleoperation
Video Teleoperation
These clips visualize SONIC ’s ability to track diverse human motions—including walking, running, turning, jumping, dancing, floor motions for teleoperation. The videos emphasize our robust zero-shot imitation of out-of-distribution motion sequences
Full-Body VR Control
Using a PICO headset, ankle trackers, and controllers, full-body pose streams drive SONIC via the human-motion encoder. The robot performs walking, sidestepping, crouching, and loco-manipulations with low latency.
3-Point VR Teleoperation
A lightweight teleoperation mode uses only the headset and handheld controllers. The user provides upper body commands via the controllers and the navigation command for the planner to generate natural lower body movements. Videos show SONIC performing mobile manipulation, navigation, reaching, and fast upper-body control. Hybrid upper-body commands with lower-body motion planning provide stable teleoperation interface.
S1.2 Interactive Motion Control
These videos demonstrate the real-time kinematic motion planner coupled with the universal tracking policy.
Interactive Locomotion
Users control the humanoid through continuous velocity and heading commands. SONIC replans every $\sim$100 ms and produces highly responsive and natural movements.
Walking Style
Clips illustrate smooth transitions between motion styles such as drunken walk, injured walk, happy walk, stealth walk, and sprinting. Our planner’s inbetweening capabilities enable seamless blends among styles.
Interactive Boxing
SONIC executes jabs, hooks, stance shifts, and reactive upper-body movements with fluidity. Unlike clip-switching methods, SONIC produces continuous, unbroken sequences with minimal artifacts.
Squatting, Kneeling, and Crawling
Videos show arbitrary pelvis-height control (0.3–0.8 m), kneeling variations, elbow crawling, hand–knee crawling, and transitions. These skills enable locomotion in confined spaces and downstream teleoperation behaviors.
S1.3 Multi-Modal Motion Generation
These videos highlight SONIC ’s universal token space by accepting diverse motion interfaces—text and music.
Text-Driven Motion.
SONIC responds to natural-language commands such as “walk forward”, “punch with your left hand”, or “act like a monkey” and can track the generated motion from GENMO (27).
Music-Conditioned Motion.
SONIC performs dance motions synchronized to tempo and musical phrasing, adapting to rhythmic and stylistic cues.
These videos demonstrate seamless transitions among modalities without training a new controller.
S1.5 Foundation-Model-Driven Loco-Manipulation
These clips show a GR00T N1.5 vision–language–action (VLA) model driving SONIC via the same teleoperation interface.
  • The VLA predicts head and wrist poses, base height, and navigation commands.
  • SONIC executes these signals through its kinematic planner and universal control decoder.
  • Videos show autonomous apple-to-plate mobile pick-and-place achieving 95% success.
This integration demonstrates SONIC as a fast, reactive System 1 whole-body controller under a system 2 VLA planner.
S2. Supplementary Tables
This section includes four tables that provide implementation details of SONIC. These tables provide complete specifications for the network architecture, training hyperparameters, reward definitions, and domain randomization settings used to obtain the large-scale motion tracking results presented in the main paper.
Table S2.1: Network Architecture
Table 1 reports the full configuration of the encoder–decoder architecture used by SONIC, including:
  • robot-motion encoder, human-motion encoder, and hybrid encoder;
  • universal token quantizer parameters (FSQ levels, token dimension);
  • robot-control decoder and auxiliary robot-motion decoder;
  • layer dimensions, MLP depth, activation functions, and latent sizes.
Table S2.2: Training Hyperparameters
Table 1 lists all training configurations, including PPO hyperparameters, rollout length, discount factors, etc.
Table S2.3: Reward Function
Table 3 provides a complete breakdown of the reward terms used for large-scale motion tracking, including:
  • root position/orientation matching,
  • joint position and velocity tracking,
  • body-link pose and velocity terms,
  • action regularization and smoothness,
  • penalty terms for joint limits and undesired contacts.
Table S2.4: Domain Randomization
Table 4 reports the ranges and distributions for all physical and kinematic randomization parameters used during training. These settings are applied during RL to ensure SONIC transfers robustly to the real Unitree G1.
Together, these four tables fully specify SONIC ’s training and implementation details to support reproducibility and further research.
Module Architecture Dims
Network configuration
Quantizer FSQ token dimensions $D_{z}$; quantization levels $L_{z}$
Encoder (g1) MLP hidden=[2048,1024,512,512]=[2048,1024,512,512]
Encoder (teleop) MLP hidden =[2048,1024,512,512]=[2048,1024,512,512]
Encoder (smpl) MLP hidden =[2048,1024,512,512]=[2048,1024,512,512]
Decoder (actions) MLP hidden =[4096,4096,2048,2048,1024,1024,512,512]=[4096,4096,2048,2048,1024,1024,512,512]
Decoder (refs) MLP hidden =[2048,1024,512,512]=[2048,1024,512,512]
Action dimension Diagonal Gaussian $29$
Critic MLP hidden =[4096,4096,2048,2048,1024,1024,512,512]=[4096,4096,2048,2048,1024,1024,512,512]
Motion command
Future frames $F_{r}=F_{h}=F_{m}=10$ frames
Frame interval $\Delta t_{r}=t_{m}=0.1s$, $\Delta t_{h}=0.02s$
Table 1: Universal control policy hyperparameters.
Training hyperparameter Value
Num parallel envs per GPU $4096$
Num steps per env 24
Learning epochs 55
Num mini-batches 44
Discount $\gamma$ $0.99$
GAE $\lambda$ $0.95$
Clip parameter $0.2$
Entropy coefficient $0.013$
Value loss coefficient $1.0$
Actor learning rate $2{\times}10^{-5}$
Critic learning rate $1{\times}10^{-3}$
Max gradient norm $0.1$
Desired KL $0.01$
Adaptive LR min/max $1{\times}10^{-5}$ , $2{\times}10^{-4}]$
Init noise std $0.05$
Actor std clamp min/max $[0.001,0.5]$
Adaptive sampling bin size 1s
Adaptive sampling failure rate cap $\beta=200$
Adaptive sampling blending hyperparameter $\alpha=0.1$
Table 2: Training hyperparameters.
Reward term Equation Weight
Tracking rewards $\mathcal{R}({\bm{s}^{\text{p}}_{t}},{\bm{s}^{\text{g}}_{t}})$
Penalty terms $\mathcal{P}({\bm{s}^{\text{p}}_{t}},\bm{a}_{t})$
Action rate $r_{\text{act}}(t)=\|\bm{a}_{t}-\bm{a}_{t-1}\|_{2}^{2}$ -0.1
Table 3: Reward design. Superscript $g$: goal/target state from $\bm{s}^{\text{g}}_{t}$}; superscript $p$: proprioceptive/current state from $\bm{s}^{\text{p}}_{t}$}; $\mathcal{B}$: set of tracked body links; superscript “rel”: quantities relative to root frame.
Domain Randomization Sampling Distribution
Physical parameters
Static friction coefficients $\mu_{s}\sim\mathcal{U}[0.3,1.6]$
Dynamic friction coefficients $\mu_{d}\sim\mathcal{U}[0.3,1.2]$
Restitution coefficient $e\sim\mathcal{U}[0,0.5]$
Default joint positions $\bm{q}_{0}\sim\bm{q}_{0}+\mathcal{U}[-0.01,0.01]$
Root velocity perturbations (external pushes)
Push duration $\Delta t\sim\mathcal{U}[1,3]\text{s}$
Target motion perturbations ($\bm{s}^{\text{g}}_{t}$})
Target joint jitter $\Delta\bm{q}_{t}^{g}\sim\mathcal{U}[-0.1,0.1]$